# Adaptive variational quantum dynamics simulations with compressed circuits and fewer measurements

### v1

@@ V Leveraging classical computation in AVQDS(T)
For practical AVQDS(T) calculations on QPUs, most quantum resource intensive component is the measurement overhead for the $M_{\mu\nu}$ elements, which scales quadratically with the number of variational parameters $N_{\boldsymbol{\theta}}$ Yao et al. (2021). Since $M_{\mu\nu}$ is equivalent to the real part of the quantum geometric tensor (QGT) or quantum Fisher information matrix Meyer (2021) and therefore has wide applications in quantum science, methods to efficiently measure it on QPUs are under active development Meyer (2021); Kolotouros et al. (2024); Liu et al. (2019); Gacon et al. (2021), primarily utilizing stochastic approaches Meyer (2021); Kolotouros et al. (2024). Here we propose a complementary approach to save quantum resources without sacrificing accuracy by maximally leveraging classical computational resources thanks to the algorithmic structure of AVQDS(T). Specifically, one can capitalize on the observation that the measurement of an element $M_{\mu\nu}$ ($\mu\leq\nu$) only invokes the ansatz circuit fragment $\hat{U}_{\nu}=\prod_{\gamma=1}^{\nu}e^{-i\theta_{\gamma}\hat{\mathcal{A}}_{\gamma}}$ up to a depth defined by the $\nu$th parameter, rather than the full ansatz circuit containing $N_{\boldsymbol{\theta}}$ parameters Yao et al. (2021). This implies that an inner block of the matrix $M$ can always be measured using low-depth circuits, amenable to efficient classical computation with high accuracy. In particular, tensor network (TN) based techniques can be naturally exploited to boost the efficiency of AVQDS(T) calculations. For dynamics simulations starting with easily prepared states, such as product states as demonstrated in this work, the AVQDS(T) circuit depth gradually grows from $\mathcal{O}(1)$ as time evolves. This allows the AVQDS(T) simulations for very early times to be executed accurately using only classical resources. For instance, matrix product states can efficiently simulate arbitrarily wide quantum circuits with depths up to around 10, since the increase of bond dimension $\chi$ incurred by each layer of gates is upper bounded by a factor of 2 (if the entangling gates are between two locally nearby qubits). Only when the AVQDS(T) circuit depth grows beyond $\sim\log_{2}(\chi)$, or equivalently $t>\log_{2}(\chi)$ assuming linear circuit depth growth with initial time, are QPUs required to measure the remaining $M_{\mu\nu}$ elements, which are beyond the reach of classical approaches.
@@
As a proof-of-principle demonstration of the proposed approach, we apply AVQDS(T) to quench dynamics simulations of the TFIM with $N_{q}=10$ on a simulator including shot noise. An inner block of the $M$ matrix up to a dimension set by a threshold circuit depth $d_{c}$ is accurately evaluated using a classical algorithm, namely exact diagonalization for simplicity. The remaining $M_{\mu\nu}$ elements are measured with $N_{s}$ “shots” in the manner described in Sec. IV.1, mimicking the noisy matrix elements that would be obtained on a QPU. (Recall that shot noise is incorporated as Gaussian random noise with standard deviation determined by $N_{s}$, which can also be viewed as a proxy for other types of noise one would encounter on a QPU.) Since the focus here is on the efficacy of partial measurement of the $M$ matrix, we set the $N_{\boldsymbol{\theta}}$ elements of $V$ (5) to their state vector values (i.e., those that would be obtained with $N_{s}=\infty$). We contrast the performance of AVQDS(T) with that of HVA with $L=10$ layers. The HVA circuit amounts to a depth of $30$, which is comparable to that of the final AVQDS(T) circuit. In contrast to AVQDS(T), the $M$ matrix for VQDS with HVA has a fixed dimension and only the inner block of dimension $d_{c}(<30)$ is evaluated “classically” (i.e., without noise). Therefore, shot noise impacts the entire fixed-ansatz VQDS simulation from the very beginning, unlike in AVQDS(T) where it only becomes important when the adaptive ansatz depth exceeds $d_{c}$.
@@
To demonstrate the impact of noise in AVQDS(T), $d_{c}$ needs to be less than the saturated circuit depth; we choose $d_{c}=21$ and $27$. We also set an upper bound of $30$ on the ansatz circuit depth for AVQDS(T), matching that of the fixed-depth HVA circuit. In Fig. 6 (c)-(f), we plot the wavefunction infidelity as a function of time for the noisy simulations. The time traces are averaged over 20 independent runs for each parameter set, with the error bars showing the standard deviation. The light purple regions denote the acceptance range with $f>0.90$. AVQDS(T) simulations are noiseless for times corresponding to circuit depths $d\leq d_{c}$, leading to more accurate results than HVA in the early stage of the simulation, even though the AVQDS(T) circuit is much shallower. This is in contrast with the noiseless case shown in Fig. 6 (a), where the HVA wavefunction, in general, has a higher fidelity when $t$ is small. After the noise kicks in when $d>d_{c}$, the AVQDS(T) wavefunction infidelity increases more rapidly, creating a kink in the trajectory. This demonstrates that noise has a significant detrimental effect on the dynamics simulation fidelity even when only a small fraction of the components of $M$ are noisy. For both $d_{c}=21$ and $27$, the wavefunctions generated by AVQDS(T) remain more accurate than HVA for a certain time period. Moreover, when $d_{c}$ is increased to $27$, simulations with AVQDS (T) appear to stay in the acceptance range noticeably longer than HVA, as shown in Fig. 6 (e) and (f). While the time traces generated by the two methods eventually merge, these results demonstrate that with the ability to leverage classical algorithms, dynamical simulations with AVQDS(T) can achieve more accurate results for a considerable time duration by deferring the onset of noise. Even more importantly, by adaptively expanding the ansatz over time and leveraging classical simulations until $d>d_{c}$, we shift the use of quantum resources to where they are most needed. Quantum resources are only used beyond classical simulation times and only to simulate those circuits that are cannot be executed on classical computers. For example, for $d_{c}=21$ and $N_{s}=10^{4}$, while HVA and AVQDS(T) stays in the acceptance range for about the same time ($t<2.5$) as shown in Fig. 6 (c), the total number of measurements required in AVQDS(T) is only $\sim 10\%$ of that in HVA for this time duration.
Finally, for reference, VQDS results with HVA obtained exclusively on noisy simulators, i.e., without implementing the hybrid classical/quantum strategy by setting $d_{c}=0$, are also included in Fig. 6 (c)-(f), where one can see that even for a fixed ansatz, the dynamical simulations can still benefit greatly from accurately evaluating an inner section of the $M$ matrix using a classical algorithm.
@@ VI Conclusion
We report progress on enabling adaptive variational quantum dynamics simulations on noisy quantum hardware with the goal to reach beyond-classical simulation times. Specifically, we discuss three improvements compared to the original AVQDS method Yao et al. (2021): (i) we implement a TETRIS approach to significantly compress the circuit depth and to reduce the number of two-qubit entangling gates. Unlike the original AVQDS, in which the operators are added to the variational ansatz one at a time, AVQDS(T) adds a series of operators acting on disjoint sets of qubits at each iteration. (ii) We benchmark a noise-resilient scheme to solve the dynamical equations of motion for the variational parameters based on a truncation of the eigenvalues of the matrix $M$ (the real part of the QGT), and (iii) we leverage classical computing resources by computing a part of $M$ on classical hardware and use quantum resources only during beyond-classical timescales and circuit depths. We believe that this synergistic interplay between classical and quantum computing is one of the main advantages of variational approaches to quantum dynamics simulations, where one can defer the use of quantum resources to the beyond-classical regime and employ them when they are most effective.
@@
To benchmark these advancements, we demonstrated AVQDS(T) calculations for the TFIM, MFIM and HM with $N_{q}\geq 10$. In addition to greatly reducing the circuit depth compared with AVQDS and Trotter decomposition, AVQDS(T) also generates circuits with fewer 2-qubit unitaries, better suited for current and near-term QPUs. The circuit compression is relevant to reach longer simulation times as the circuit depth sets the execution time on hardware, which is upper bounded by the device’s coherence time. We also compared several methods for solving the equations of motion controlling the variational quantum dynamics. While the bounded least-square minimization, Tikhonov regularization, and truncation methods all work satisfactorily on noiseless simulators, the truncation method is found to be more noise-resilient as demonstrated by simulations in the presence of shot noise. Finally, we propose a way to save quantum resources in AVQDS(T) calculations by exploiting classical algorithms to reduce the quantum measurement overhead of the quantum Fisher information matrix $M$, which is the most demanding part of the algorithm. Classical computation can be leveraged in the whole simulation period because an inner block of $M$ involving low-depth ansatz circuit fragments can always be efficiently evaluated. The gradual circuit growth entailed by the AVQDS(T) algorithm also renders it feasible to simulate the dynamics at very early times using only classical computations. With a proof-of-principle demonstration including shot noise effects, we show that AVQDS(T) aided with classical evaluations of a sub-block of $M$ produces more accurate dynamics than HVA for a notable time duration at comparable circuit depths. In view of the fact that classical algorithms like TN-based approaches are being actively developed for quantum circuit simulations Ayral et al. (2023); Tindall et al. (2024), we envision this synergistic engagement of both quantum and classical resources, in particular the partial classical evaluation of the $M$-matrix, will help boost AVQDS(T) simulations to large-size systems for showcasing quantum utility Herrmann et al. (2023); Kim et al. (2023).
@@ Acknowledgements
Appendix A Solubility of Eq. (2) for a singular matrix $M$

### v2

@@ V Leveraging classical computation in AVQDS(T)
For practical AVQDS(T) calculations on QPUs, most quantum resource intensive component is the measurement overhead for the $M_{\mu\nu}$ elements, which scales quadratically with the number of variational parameters $N_{\boldsymbol{\theta}}$ [34]. Since $M_{\mu\nu}$ is equivalent to the real part of the quantum geometric tensor (QGT) or quantum Fisher information matrix [56] and therefore has wide applications in quantum science, methods to efficiently measure it on QPUs are under active development [56, 57, 58, 59], primarily utilizing stochastic approaches [56, 57]. Here we propose a complementary approach to save quantum resources without sacrificing accuracy by maximally leveraging classical computational resources thanks to the algorithmic structure of AVQDS(T). Specifically, one can capitalize on the observation that the measurement of an element $M_{\mu\nu}$ ($\mu\leq\nu$) only invokes the ansatz circuit fragment $\hat{U}_{\nu}=\prod_{\gamma=1}^{\nu}e^{-i\theta_{\gamma}\hat{\mathcal{A}}_{\gamma}}$ up to a depth defined by the $\nu$th parameter, rather than the full ansatz circuit containing $N_{\boldsymbol{\theta}}$ parameters [34]. This implies that an inner block of the matrix $M$ can always be measured using low-depth circuits, amenable to efficient classical computation with high accuracy. In particular, tensor network based techniques can be naturally exploited to boost the efficiency of AVQDS(T) calculations. For dynamics simulations starting with easily prepared states, such as product states as demonstrated in this work, the AVQDS(T) circuit depth gradually grows from $\mathcal{O}(1)$ as time evolves. This allows the AVQDS(T) simulations for very early times to be executed accurately using only classical resources. For instance, matrix product states can efficiently simulate quantum circuits of arbitrary width with depths up to around 10, assuming the entangling gates are applied between neighboring qubits, such as in one-dimensional local spin models. This limitation arises because the bond dimension, $\chi$, increases with each layer of local entangling gates, with the growth being upper-bounded by a factor of 2. QPUs become necessary only when the AVQDS(T) circuit depth surpasses $\sim\log_{2}(\chi)$ at a certain simulation time $t$. Beyond this threshold, QPUs are required to measure the remaining $M_{\mu\nu}$ elements, which lie beyond the capabilities of classical methods.
@@
As a proof-of-principle demonstration of the proposed approach, we apply AVQDS(T) to quench dynamics simulations of the TFIM with $N_{q}=10$ on a simulator including shot noise. An inner block of the $M$ matrix up to a dimension set by a threshold circuit depth $d_{c}$ is accurately evaluated using a classical algorithm, namely exact diagonalization for simplicity. The remaining $M_{\mu\nu}$ elements are measured with $N_{s}$ “shots” in the manner described in Appendix B, mimicking the noisy matrix elements that would be obtained on a QPU. (Recall that shot noise is incorporated as Gaussian random noise with standard deviation determined by $N_{s}$, which can also be viewed as a proxy for other types of noise one would encounter on a QPU.) Since the focus here is on the efficacy of partial measurement of the $M$ matrix, we set the $N_{\boldsymbol{\theta}}$ elements of $V$ (5) to their state vector values (i.e., those that would be obtained with $N_{s}=\infty$). We contrast the performance of AVQDS(T) with that of HVA with $L=10$ layers. The HVA circuit amounts to a depth of $30$, which is comparable to that of the final AVQDS(T) circuit. In contrast to AVQDS(T), the $M$ matrix for VQDS with HVA has a fixed dimension and only the inner block of dimension $d_{c}(<30)$ is evaluated “classically” (i.e., without noise). Therefore, shot noise impacts the entire fixed-ansatz VQDS simulation from the very beginning, unlike in AVQDS(T) where it only becomes important when the adaptive ansatz depth exceeds $d_{c}$.
@@
To demonstrate the impact of noise in AVQDS(T), $d_{c}$ needs to be less than the saturated circuit depth; we choose $d_{c}=21$ and $27$. We also set an upper bound of $30$ on the ansatz circuit depth for AVQDS(T), matching that of the fixed-depth HVA circuit. In Fig. 5 (c)-(f), we plot the wavefunction infidelity as a function of time for the noisy simulations. The time traces are averaged over 20 independent runs for each parameter set, with the error bars showing the standard deviation. The light purple regions denote the acceptance range with $f>0.90$. AVQDS(T) simulations are noiseless for times corresponding to circuit depths $d\leq d_{c}$, leading to more accurate results than HVA in the early stage of the simulation, even though the AVQDS(T) circuit is much shallower. This is in contrast with the noiseless case shown in Fig. 5 (a), where the HVA wavefunction, in general, has a higher fidelity when $t$ is small. After the noise kicks in when $d>d_{c}$, the AVQDS(T) wavefunction infidelity increases more rapidly, creating a kink in the trajectory. This demonstrates that noise has a significant detrimental effect on the dynamics simulation fidelity even when only a small fraction of the components of $M$ are noisy. For both $d_{c}=21$ and $27$, the wavefunctions generated by AVQDS(T) remain more accurate than HVA for a certain time period. Moreover, when $d_{c}$ is increased to $27$, simulations with AVQDS (T) appear to stay in the acceptance range noticeably longer than HVA, as shown in Fig. 5 (e) and (f). While the time traces generated by the two methods eventually merge, these results demonstrate that with the ability to leverage classical algorithms, dynamical simulations with AVQDS(T) can achieve more accurate results for a considerable time duration by deferring the onset of noise. Even more importantly, by adaptively expanding the ansatz over time and leveraging classical simulations until $d>d_{c}$, we shift the use of quantum resources to where they are most needed. Quantum resources are only used beyond classical simulation times and only to simulate those circuits that are cannot be executed on classical computers. For reference, VQDS results with HVA obtained exclusively on noisy simulators by setting $d_{c}=0$, are also included in Fig. 5 (c)-(f), where one can see that even for a fixed ansatz, the dynamical simulations can still benefit greatly from accurately evaluating an inner section of the $M$ matrix using a classical algorithm.
Finally, we compare the number of distinct circuits required for quantum measurements to demonstrate the effectiveness of the hybrid classical/quantum strategy in saving quantum resources. Using $d_{c}=21$ and $N_{s}=10^{4}$ as an example [see Fig. 5(c)], for $t<2.7$ when the fidelity is within the acceptance range, the required total number of distinct circuits for AVQDS is $2.0\times 10^{6}$, while the number of distinct circuits for HVA with the same circuit-depth cutoff during the same period of time is nearly an order of magnitude larger: $1.5\times 10^{7}$. Furthermore, if HVA is run exclusively on noisy simulators ($d_{c}=0$), the number of distinct circuits is $3.6\times 10^{7}$, which is more than doubled compared with HVA that implements classical algorithms for $d\leq 21$.
@@ VI Conclusion
We report progress on enabling adaptive variational quantum dynamics simulations on noisy quantum hardware with the goal to reach beyond-classical simulation times. Specifically, we discuss three improvements compared to the original AVQDS method [34]: (i) we implement a TETRIS approach to significantly compress the circuit depth. Unlike the original AVQDS, in which the operators are added to the variational ansatz one at a time, AVQDS(T) adds a series of operators acting on disjoint sets of qubits at each iteration. (ii) We benchmark a noise-resilient scheme to solve the dynamical equations of motion for the variational parameters based on a truncation of the eigenvalues of the matrix $M$ (the real part of the QGT), and (iii) we leverage classical computing resources by computing a part of $M$ on classical hardware and use quantum resources only during beyond-classical timescales and circuit depths. We believe that this synergistic interplay between classical and quantum computing is one of the main advantages of variational approaches to quantum dynamics simulations, where one can defer the use of quantum resources to the beyond-classical regime and employ them when they are most effective.
@@
To benchmark these advancements, we demonstrated AVQDS(T) calculations for the TFIM, MFIM and HM with $N_{q}\geq 10$. In addition to greatly reducing the circuit depth compared with AVQDS and Trotter decomposition, AVQDS(T) also generates circuits with fewer 2-qubit unitaries, better suited for current and near-term QPUs. The circuit compression is relevant to reach longer simulation times as the circuit depth sets the execution time on hardware, which is upper bounded by the device’s coherence time. In a comparative study, we demonstrate that the required quantum resources are comparable in AVQDS(T) and an alternative adaptive pVQD method in order to achieve similar accuracies, while the overall circuit depth in AVQDS(T) is significantly shallower than the latter. Simulations implementing realistic noise channels show that under a typical error rate of $10^{-4}$ for single-qubit gates, the quantum observables agree reasonably well with exact values for small error rates for two-qubit gates ($10^{-4}$ and $10^{-3}$), while showing qualitative differences with exact values for larger two-qubit error rates. Finally, we propose a way to save quantum resources in AVQDS(T) calculations by exploiting classical algorithms to reduce the quantum measurement overhead of the quantum Fisher information matrix $M$, which is the most demanding part of the algorithm. Classical computation can be leveraged in the whole simulation period because an inner block of $M$ involving low-depth ansatz circuit fragments can always be efficiently evaluated. The gradual circuit growth entailed by the AVQDS(T) algorithm also renders it feasible to simulate the dynamics at very early times using only classical computations. With a proof-of-principle demonstration including shot noise effects, we show that AVQDS(T) aided with classical evaluations of a sub-block of $M$ produces more accurate dynamics than HVA for a notable time duration at comparable circuit depths. In view of the fact that classical algorithms like TN-based approaches, both rigorous and approximate, are being actively developed for quantum circuit simulations [60, 61], we envision this synergistic engagement of both quantum and classical resources, in particular the partial classical evaluation of the $M$-matrix, will help boost AVQDS(T) simulations to large-size systems for showcasing quantum utility [62, 63].
@@ Acknowledgements
Appendix A Solubility of the Equations of Motion
@@
Appendix B Technical approaches for solving the equations of motion
An essential process in both VQDS and AVQDS is to solve the equations of motion Eq. (2), which can face a numerical complication due to the singularity of the matrix $M$ as defined in Eq. (4); that is, there exist non-zero vectors $x$ that satisfy $Mx=0$. In Appendix A, we show that Eq. (2) remains solvable since the vector $V$ given in Eq. (5) is always orthogonal to the null space of $M$. In this scenario, Eq. (2) has infinitely many solutions that lie on a hyperplane in the $N_{\boldsymbol{\theta}}$-dimensional parameter space. We show in Fig. 6 (a) a schematic of the hyperplane in a 2D space, which reduces to a straight line. Obviously one cannot solve Eq. (2) by directly inverting $M$. Instead, one could try to minimize $\norm{M\dot{\boldsymbol{\theta}}-V}^{2}$ using least-square (LSQ) techniques [64]. An unbounded LSQ search can in principle reach any solution on the hyperplane, which can have a large magnitude. Consequently, a tiny time increment $\Delta t$ must be taken so that the changes in the rotation angles $\Delta\theta_{\mu}=\dot{\theta}_{\mu}\Delta t$ remain small enough to guarantee a continuous integration of the equations of motion. A remedy for this problem is to set a bound $b$ for the LSQ method, so the search is restricted to be within a hypercube of side $2b$ centered at the origin. An exact solution can still be found as long as the hypercube intersects with the hyperplane [see Fig. 6 (a)].
Tikhonov regularization [65] is another popular method for solving linear systems of equations of the form (2) with singular or ill-conditioned $M$. In this method, the solution to Eq. (2) is computed as $\dot{\boldsymbol{\theta}}=(M+\varepsilon I)^{-1}V$, where $\varepsilon$ is a small number and $I$ is the identity matrix. To better understand the effect of Tikhonov regularization, we first diagonalize the symmetric matrix $M$ with a unitary transformation $M=U\Lambda U^{\dagger}$, where $\Lambda$ is a diagonal matrix containing the eigenvalues $\lambda_{\mu}$ of $M$ in ascending order and where each column of $U$ gives the corresponding eigenvector. The eigenvalues $\lambda_{\mu}\geq 0$ since $M$ is positive semi-definite (see Appendix A). Assume the null space of $M$ has dimension $N_{\text{null}}$; then, $\lambda_{\mu}=0$ for $1\leq\mu\leq N_{\text{null}}$, and the null space of $M$ is spanned by the first $N_{\text{null}}$ eigenvectors of $M$: $\text{Null}(M)=\text{Span}\{u_{1},u_{2},\cdots,u_{N_{\text{null}}}\}$ where $u_{\mu}$ is the $\mu^{\text{th}}$ column of $U$. Under Tikhonov regularization, $\dot{\boldsymbol{\theta}}=(M+\varepsilon I)^{-1}V=U(\Lambda+\varepsilon I)^{-1}U^{\dagger}V$. $(\Lambda+\varepsilon I)^{-1}$ is also a diagonal matrix dominated by the first $N_{\text{null}}$ diagonal elements ($\varepsilon^{-1}$). However, since $V\perp\text{Null}(M)$, the first $N_{\text{null}}$ elements of $U^{\dagger}V$ are zero. Thus, as long as $\varepsilon$ remains large compared to the machine accuracy, the first $N_{\text{null}}$ elements of the resulting vector $(\Lambda+\varepsilon I)^{-1}U^{\dagger}V$, which gives the projection of $\dot{\boldsymbol{\theta}}$ onto $\text{Null}(M)$, are essentially zero. Meanwhile, the Tikhonov parameter $\varepsilon$ should also be much smaller than the nonzero eigenvalues $\lambda_{\mu}>0$. Thus, $\dot{\boldsymbol{\theta}}$ calculated in this way is close to the unique solution of Eq. (2) that is orthogonal to $\text{Null}(M)$. Such a solution is also schematically shown in Fig. 6 (a).
Alternatively, the coefficient matrix $M$ can be regularized by performing a truncation on its eigenvalues [43]. After diagonalizing $M$, Eq. (2) can be rewritten as $\Lambda(U^{\dagger}\dot{\boldsymbol{\theta}})=U^{\dagger}V$. Since the first $N_{\text{null}}$ elements of both $\lambda_{\mu}$ and $U^{\dagger}V$ are zero, a solution of the above linear system of equations can be straightforwardly obtained by letting $(U^{\dagger}\dot{\boldsymbol{\theta}})_{\mu}=0$ for $1\leq\mu\leq N_{\text{null}}$, and $(U^{\dagger}\dot{\boldsymbol{\theta}})_{\mu}=(U^{\dagger}V)_{\mu}/\lambda_{\mu}$ for $N_{\text{null}}<\mu\leq N_{\boldsymbol{\theta}}$. In other words, $U^{\dagger}\dot{\boldsymbol{\theta}}$ is an $N_{\boldsymbol{\theta}}$-dimensional vector whose first $N_{\text{null}}$ elements are zeros. After $U^{\dagger}\dot{\boldsymbol{\theta}}$ is obtained, $\dot{\boldsymbol{\theta}}$ is readily available as $\dot{\boldsymbol{\theta}}=U(U^{\dagger}\dot{\boldsymbol{\theta}})$. This procedure again leads to the unique solution located in the subspace orthogonal to $\text{Null}(M)$ as shown in Fig. 6 (a). In practice, $N_{\text{null}}$ is identified by truncating the eigenvalues based on a threshold value $\varepsilon$; in other words, $N_{\text{null}}=\left|\{\lambda_{\mu}|\lambda_{\mu}\leq\varepsilon\}\right|$ where ||$\cdot$|| ⋅ | denotes the size of a set. On noiseless simulators, this method can be interpreted as a modified Tikhonov regularization, where the uniform shift of the eigenvalues of $M$ by $\varepsilon I$ is replaced with a partial shift involving only the negligible eigenvalues within the null space. However, when noise is present, the first $N_{\text{null}}$ eigenvalues of $M=M_{0}+\delta M$ will take nonzero values $\delta\lambda_{\mu}$. Here, $M_{0}$ and $\delta M$ denote the exact coefficient matrix and the noise, respectively. At the same time, $u_{\mu}^{\dagger}V$ ($\mu\leq N_{\text{null}}$) will not vanish either, and their values $\delta(u_{\mu}^{\dagger}V)$ are also determined by the noise. Consequently, the projection of $\dot{\boldsymbol{\theta}}$ onto $\text{Null}(M_{0})$, calculated within Tikhonov regularization via $(\delta\lambda_{i}+\varepsilon)^{-1}\delta(u_{i}^{\dagger}V)$, can be large because it involves division of small quantities. On the other hand, such operations are largely avoided in the truncation method by effectively truncating the eigenvalues $\delta\lambda_{\mu}$ that are only nonzero due to noise, resulting in a solution in the vicinity of the orthogonal subspace of $\text{Null}(M_{0})$. This method, which we call the “truncation” method, has been applied to variational quantum imaginary time evolution [44]. Here, we examine its performance on simulating real-time quantum dynamics on both noiseless and noisy simulators.
Figure 6: (a) Schematic showing different approaches for solving Eq. (2) in a 2D parameter space. The red dashed line shows the null space of $M$ (the collection of all $x$ satisfying $Mx=0$). The solid black line parallel to $\text{Null}(M)$ is the hyperplane containing all the solutions of Eq. (2). The black circles show the solutions that can be obtained using different methods. In particular, the Tikhonov regularization and the truncation methods can locate the unique solution orthogonal to $\text{Null}(M)$. (b) The growth of the number of unitaries $N_{\boldsymbol{\theta}}$ as a function of time, (c) the time evolution of the wavefunction infidelity, and (d) the time increment $\Delta t=\Delta\theta_{m}/\max_{\mu}\absolutevalue{\dot{\theta}_{\mu}}$ | at each time step as a function of time, during the quench dynamics simulation of the TFIM with $N_{q}=12$, using the methods outlined in (a) to solve Eq. (2).
We compare the above-mentioned methods for solving the equations of motion Eq. (2) by performing AVQDS(T) on the TFIM with $N_{q}=12$. We here use method 3 of Fig. 1 to construct the compressed circuits. $\varepsilon$ is set to $10^{-6}$ for both the Tikhonov regularization and truncation methods. For the bounded LSQ search, we set the bound $b=5$. The variational parameters $\boldsymbol{\theta}$ are updated according to the Euler method $\Delta\boldsymbol{\theta}=\dot{\boldsymbol{\theta}}\Delta t$ with $\Delta t=\Delta\theta_{m}/\max_{\mu}\absolutevalue{\dot{\theta}_{\mu}}$ | [49], where the maximal allowed change in rotation angles $\Delta\theta_{m}$ is set to 0.005. Fig. 6 (b) plots the growth of the number of unitaries $N_{\boldsymbol{\theta}}$ as a function of $t$, which shows that the different methods produce circuits of similar complexity. We show the infidelity $1-f$ of |$\ket{\Psi[\boldsymbol{\theta}(t)]}$ and the time increment at each time step ($\Delta t$) as a function of $t$ in Fig. 6 (c) and (d), respectively. The fidelity $f$ is defined as $f\equiv\absolutevalue{\langle\Psi[\boldsymbol{\theta}(t)]|\,\Psi_{\text{exact}}(t)\rangle}^{2}$, where |$\ket{\Psi_{\text{exact}}(t)}=e^{-i\hat{\mathcal{H}}t}\ket{\Psi_{0}}$ is obtained by directly applying the time-evolution operator. All the curves coincide for $t<2$ in both Fig. 6 (c) and (d), indicating the matrix $M$ is well-conditioned, and all the methods can effectively locate the unique solution to Eq. (2) during this time period.
In the current NISQ era, it is important to examine the noise resilience of different strategies for solving Eq. (2) in a noisy environment. Here, we only consider the shot noise on the matrix $M$—resulting from a fixed set of circuit samples or “shots” performed per matrix element—to single out the effect of these solvers on regularizing $M$. The matrix elements of $M$ can be obtained by calculating the probability $p$ that an ancilla qubit measurement yields 1 [34], such that $M_{\mu\nu}=(2p-1)/4$. The variance $\sigma_{M_{\mu\nu}}^{2}=p(p-1)/(4N_{s})$ where $N_{s}$ is the number of shots used to measure $M_{\mu\nu}$. We assume that all matrix elements are measured with the same number of shots. In our calculations, we simulate the shot noise by replacing the noiseless value $M_{\mu\nu}$ with a random number drawn from a Gaussian distribution with mean $M_{\mu\nu}$ and variance $\sigma_{M_{\mu\nu}}^{2}$.
Since the bounded LSQ method requires a predetermined bound that is generally problem-specific, we focus on the Tikhonov regularization and truncation methods, taking the TFIM with $N_{q}=10$ as an example. A fixed HVA is adopted to ensure that the quantum circuits are the same throughout the simulation for both methods. Relatively deeper circuits are required in the presence of noise; thus, we choose $L=15$ and 20. $N_{s}$ is set to moderate values $10^{4}$ and $10^{5}$. The parameter $\varepsilon$ controls the small positive value added to the diagonal of the matrix $M$ in the Tikhonov regularization method, or the threshold for truncating the eigenvalues of $M$ in the truncation method. Relative to noiseless simulations, a larger $\varepsilon$ is needed so as to avoid being overwhelmed by the noise. On the other hand, if $\varepsilon$ is too large, it causes too much distortion to the matrix $M$. After some pre-experimentation, we choose $\varepsilon=10^{-2}$ and $10^{-3}$ for the Tikhonov regularization method, and $\varepsilon=10^{-3}$ and $10^{-4}$ for the truncation method. These parameters yield the same order of magnitude for the standard error of $M_{\mu\nu}$ given $N_{s}=10^{4}$ and $10^{5}$. We implement a fixed time step $\Delta t$ for this comparison so that the total computation cost is the same for both methods. $\Delta t$ is set to 0.002 for $L=15$ and 0.005 for deeper circuits with $L=20$. 20 independent runs were performed for each parameter set. Figure 7 gives the time traces of the wavefunction infidelity, in which the region shaded in light purple shows the “acceptance range” with $f>0.9$. The contrast is clear: the simulations using Tikhonov regularization escape the acceptance range shortly after being launched, while those performed with the truncation method stay within the acceptance range for a much longer time, which increases with the number of HVA layers or the number of shots. The failure of the Tikhonov regularization method is due to its difficulty in curbing the components of $\dot{\boldsymbol{\theta}}$ on the null space of the unperturbed coefficient matrix, resulting in large $|\dot{\boldsymbol{\theta}}|$. One has to reduce $\Delta t$ by about an order of magnitude in order to make it work, greatly increasing the required quantum resources. The truncation method with $\varepsilon=10^{-3}$ works the best for all the parameter sets that have been examined, and will be used in subsequent noisy simulations. With this optimal setting, the average number of distinct circuits for the duration within the acceptance range is about $4\times 10^{11}$ using $N_{s}=10^{4}$ for both $L=15$ and $L=20$. Note that a smaller $\Delta t$ is used for $L=20$. This number increases to $10^{12}$ for $N_{s}=10^{5}$ due to the longer duration within the acceptance range. Quantum resources at this level remain challenging in the NISQ era, and manageable jobs on real devices typically feature smaller system sizes or shallower ansätze.